# Towards Galactic Archaeology with Inferred Ages of Giant Stars From Gaia Spectra

### v1

@@ 6 Summary
To illustrate the utility of SIDRA-XP, we apply the trained model to a large sample of 2,218,154 stars from Gaia’s XP stellar parameter data, selecting stars that meet specific accuracy criteria from Fallows & Sanders (2024). Our analysis mapped the chronological and chemical evolution of the Galactic disc, revealing distinct phases in the disc’s history: an early metal-poor phase, a rapid increase in metallicity likely linked to the GSE merger, and the ongoing formation of the thin disc. The results provide a hint of a gas-rich merger event around 7 Gyr ago, potentially associated with the first infall of the Sagittarius dwarf galaxy, as indicated by a sequence of stars with lower [Fe/H] and higher $[\mathrm{\alpha/M}]$.
@@ Appendix B Model trained with the Gaia Gsp-Spec stellar parameters
The Gaia DR3 provides stellar parameters derived from the RVS spectra (GSP-Spec, Recio-Blanco et al., 2023). We cross-match the BINGO data with Gaia DR3 stars that have GSP-Spec measurements for the effective temperature ($T_{\mathrm{eff,GSP-Spec}}$), surface gravity ($\log g_{\mathrm{GSP-Spec}}$), and metallicity ($[\mathrm{M/H}]_{\mathrm{GSP-Spec}}$), resulting in a dataset of 6,924 stars, referred to as the BINGO-GSP-Spec dataset. An artificial neural network (ANN) model is trained using 80% of these cross-matched data, with $T_{\mathrm{eff,GSP-Spec}}$, $\log g_{\mathrm{GSP-Spec}}$ and $[\mathrm{M/H}]_{\mathrm{GSP-Spec}}$ as the input features and BINGO age as the output. We apply the same data augmentation techniques to the training data as described in Sec. 3.1 and optimise the hyperparameters using Optuna. The predicted age for the unseen test data—i.e., the 20% of the original dataset not used for training—is shown in the left panel of Fig. 11. The ANN model trained on GSP-Spec parameters shows significant residuals, indicating discrepancies between the predicted and true age estimates ($\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ [Gyr])). This suggests that the model struggles to differentiate between younger and older stars, with the residuals showing considerable spread (mean of approximately 0.20 dex) and uncertainty (standard deviation of $\sim 0.18$ dex around $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ [Gyr]) = 1).

### v2

@@ 6 Summary
To illustrate the utility of SIDRA-XP, we apply the trained model to a large sample of 2,218,154 stars from Gaia’s XP stellar parameter data, selecting stars that meet specific accuracy criteria from Fallows and Sanders (2024). Our analysis mapped the chronological and chemical evolution of the Galactic disc, revealing distinct phases in the disc’s history: an early metal-poor phase, a rapid increase in metallicity likely linked to the GSE merger, and the ongoing formation of the thin disc. The results provide a hint of a gas-rich merger event around 7 Gyr ago, potentially associated with the first infall of the Sagittarius dwarf galaxy, as indicated by a sequence of stars with lower [Fe/H] and higher $\mathrm{\alpha/M}]$.
@@ Appendix B Model trained with the Gaia Gsp-Spec stellar parameters
The Gaia DR3 provides stellar parameters derived from the RVS spectra (GSP-Spec, Recio-Blanco et al., 2023). We cross-match the BINGO data with Gaia DR3 stars that have GSP-Spec measurements for the effective temperature ($T_{\mathrm{eff,GSP-Spec}}$), surface gravity ($\log g_{\mathrm{GSP-Spec}}$), and metallicity ($\mathrm{M/H}]_{\mathrm{GSP-Spec}}$), resulting in a dataset of 6,924 stars, referred to as the BINGO-GSP-Spec dataset. An artificial neural network (ANN) model is trained using 80% of these cross-matched data, with $T_{\mathrm{eff,GSP-Spec}}$, $\log g_{\mathrm{GSP-Spec}}$ and $\mathrm{M/H}]_{\mathrm{GSP-Spec}}$ as the input features and BINGO age as the output. We apply the same data augmentation techniques to the training data as described in Sec. 3.1 and optimise the hyperparameters using Optuna. The predicted age for the unseen test data—i.e., the 20% of the original dataset not used for training—is shown in the left panel of Fig. 11. The ANN model trained on GSP-Spec parameters shows significant residuals, indicating discrepancies between the predicted and true age estimates ($\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ [Gyr])). This suggests that the model struggles to differentiate between younger and older stars, with the residuals showing considerable spread (mean of approximately 0.20 dex) and uncertainty (standard deviation of $\sim 0.18$ dex around $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ [Gyr]) = 1).
@@
Appendix C Performance Assessment of SIDRA-XP Models with APOKASC3 Data
Figure 12: Age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA_{M21\times XP}}}$ $\mathrm{[Gyr]})$, from FS24 XP stellar parameters trained on age labels from Miglio et al. (2021) versus the target APOKASC3 age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{APOKASC3}}$ $\mathrm{[Gyr]})$ (left panel). The prediction of the age for the same sample as the left panel, but trained with SIDRA-XP using the BINGO Ciucă et al. (2021) age labels are shown in the right panel. The results are colour-coded by [Fe/H] parameter from APOGEE DR17. The upper panels show the predictions versus the target and the black dashed line indicates the identity line. The lower panels represent the residuals between the predicted age and the target age, i.e. $\log_{10}(\tau_{\mathrm{SIDRA}}/\tau_{\mathrm{APOKASC3}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at different $\mathrm{log}_{10}(\tau_{\mathrm{APOKASC3}}$ $\mathrm{[Gyr]})$ bins.
This appendix presents a comparative evaluation of two independently trained versions of the SIDRA model, as the model introduced in Section 4. Both models utilise the XP stellar parameters from Gaia DR3, as analysed in FS24, but differ in the training datasets for age calibrations. One model, referred to here as SIDRA-M21xXP, was trained using the asteroseismic ages published by Miglio et al. (2021), with the training set restricted to include stars identified as RGB stars or RC stars with masses exceeding 1.8 $\mathrm{M}_{\odot}$. The other model, SIDRA-XP, was trained using the more extensive BINGO dataset from Ciucă et al. (2021), which itself is calibrated on the ages from Miglio et al. (2021) as described in Section 4.1.
Although the Miglio et al. (2021) sample is smaller than that of BINGO, this comparison allows us to test the consistency and robustness of SIDRA predictions under different training regimes. We aim to assess the trade-off between a pristine but smaller number of the training data (2695 stars) of Miglio et al. (2021) sample and a larger training set (12,500 stars) produced with BINGO. To ensure consistency, SIDRA-M21xXP model was also constructed following the same methodology outlined in Section 4.1.
To evaluate their predictive performance, we apply both models to a subset of the APOGEE–Kepler Asteroseismic Science Consortium Catalog, version 3 (APOKASC3; Pinsonneault et al., 2025), ensuring no overlap with either training set from SIDRA-M21xXP and SIDRA-XP, which leads to 970 stars. The testing sample was constructed by selecting all stars identified as RGB, and for which reliable asteroseismic age estimates were available in APOKASC3. For each selected star, we adopt the APOKASC3_AGE_RGB.
Figure 13: Comparison between $\log_{10}$ ages from (APOKASC3; Pinsonneault et al., 2025) and those from Miglio et al. (2021), for RGB and high-mass RC stars colour-coded by APOGEE [Fe/H]DR17. The black dashed line indicates the identity line.
Comparisons between the model-predicted ages and the APOKASC3 values reveal that both models achieve comparable performance as shown in Fig. 12. Both models infer the systemically older ages than the APOKASC3 ages. This is because the asteroseismic age of Miglio et al. (2021) is systematically older than the APOKASC3 age as shown in Fig. 13 Interestingly, the SIDRA model trained on FS24 XP and BINGO data exhibited slightly improved accuracy and reduced residual scatter in the overall sample, i.e. the standard deviation of the residuals is around 0.14 dex for $\mathrm{log}_{10}(\tau_{\mathrm{APOKASC3}}$ $\mathrm{[Gyr]})=1$ for SIDRA-XP and 0.15 for $\mathrm{log}_{10}(\tau_{\mathrm{APOKASC3}}$ $\mathrm{[Gyr]})=1$ for SIDRA-M21xXP. We consider that this difference is not significant, and rather indicate that both training data provide a similar quality of the trained model.
While conducting this analysis, we observed that both models showed discrepancies for stars with APOKASC3 ages below 1 Gyr. We find that these APOKASC3 stars have high [C/N], but with a high mass. Hence, APOKASC3 inferes a younger age, but our model infers an older age. Fig. 13 shows that Miglio et al. (2021) training set includes very few RGB stars in this young age regime. Consequently, the model trained on that dataset struggles to recover accurate ages for APOKASC3 young stars, highlighting the challenge of reliable extrapolation in this regime. Although their ages may be overestimated and may contaminate older populations, their rarity suggests that the overall impact on population-level inferences is likely limited. However, this caveat should be acknowledged and kept in mind when interpreting the results.
Figure 14: Comparison between the stellar ages predicted by two independently trained models: SIDRA-M21xXP and SIDRA-XP. Both models use Gaia DR3 XP parameters from FS24 as input, but differ in their age labels: Miglio et al. (2021) for SIDRA-M21xXP and the BINGO dataset from Ciucă et al. (2021) for SIDRA-XP. The results are colour-coded by [Fe/H] parameter from APOGEE DR17. The top panel shows a 1:1 comparison of the models’ respective age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA_{M21\times XP}}}\,\mathrm{[Gyr]})$ and $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA_{XP}}}\,\mathrm{[Gyr]})$, with the black dashed line indicating the identity line. The lower panel shows the residuals between the two models, i.e. $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA_{M21\times XP}}}/\tau_{\mathrm{SIDRA_{XP}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at different $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA_{XP}}})$ bins.
Fig. 14 illustrates the close agreement between the stellar age predictions from the SIDRA-M21xXP and SIDRA-XP models, with residuals tightly centred around zero and consistent performance across the full age and metallicity range. These results indicate that the choice of age calibration, whether using Miglio et al. (2021) or BINGO, does not substantially affect the performance of the model when using XP spectra as input. In this paper, we adopt SIDRA-XP trained on the BINGO dataset, but this plot shows that SIDRA-M21xXP yields comparably reliable predictions, supporting the robustness of the method.